OpenStudy (anonymous):

write y=1/8 x+7 in standard form using integers

jimthompson5910 (jim_thompson5910):

I'll start you off \[\Large y = \frac{1}{8}x+7\] \[\Large 8*y = 8*\left(\frac{1}{8}x+7\right)\] \[\Large 8*y = 8*\left(\frac{1}{8}x\right)+8*\left(7\right)\] \[\Large 8y = x+56\] and I'll let you finish up.

jimthompson5910 (jim_thompson5910):

I multiplied both sides by 8 to clear out the fraction.